OpenStudy (anonymous):

smbdy help me ...laplace transform {t^4 + 2t^3-1}

OpenStudy (anonymous):

so laplace transform of \[t^{n}=\frac{ n! }{ s^{n+1} }\] you can use that formula

OpenStudy (anonymous):

how to solve it

OpenStudy (anonymous):

i am still early to understand it

OpenStudy (anonymous):

ok so using that formula for t^4 we would have :

OpenStudy (anonymous):

\[laplace(t^{4})=\frac{ 4! }{ t^{4+1} }\]

OpenStudy (anonymous):

when you have laplace of sum of terms like here t^4 + 2t^3+1 because of the property of linearity of the laplace you calculate laplace of each term so you would have :

OpenStudy (anonymous):

\[L({t^{4}+2t^{3}+1})=L(t^{4})+2L(t^{3})+L(1)\] you can pull constants out of the laplace

OpenStudy (anonymous):

here the question...i am still need more help to clear it..i

OpenStudy (anonymous):

for which one u need help ?

OpenStudy (anonymous):

first qstion..n the working step how to calculate it

OpenStudy (anonymous):

ok do you have the laplace transforms table in front of you ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and i already told you the formula you need for laplace of the first one you just need to apply it ..

OpenStudy (anonymous):

i am still blur..n my math fail

OpenStudy (anonymous):

ok i will do the first one all the way to the result \[L(t^{4}+2t^{3}+1)=L(t^{4})+2L(t^{3})+L(1)\] you can pull out constants in front of the laplace like i did there with the 2 in front ot t^3 no using the formula : \[L(t^{n})=\frac{ n! }{ s^{n+1} }\] we have : \[\frac{ 4! }{ s^{4+1} }+2\frac{ 3! }{ s^{3+1} }+\frac{ 0! }{ s^{0+1} }\] so our result is : \[\frac{ 24 }{ s^{5} }+\frac{ 12 }{ s^{4} }+\frac{ 1 }{ s }\] that is our solution . note that 0!=1

OpenStudy (anonymous):

L(t4+2t3-1)

OpenStudy (anonymous):

yeah that what i wrote above is the solution of that laplace transform ^^

OpenStudy (anonymous):

how u get 24

OpenStudy (anonymous):

4!=4*3*2=24

OpenStudy (anonymous):

L(3e^2t + 2e^-4t)

OpenStudy (anonymous):

\[L(e^{at}) = \frac{ 1 }{ s-a }\] you should use tables; you'll be allowed one on your exams.

OpenStudy (anonymous):

so answer is \[\frac{ 3 }{ s-2 }+ \frac{ 2 }{ s+4 }\]

OpenStudy (anonymous):

the answer only like that?

OpenStudy (anonymous):

even if ur not alowed its just few formulas not hard to remember , as i can see most of ur laplace are just table transforms

OpenStudy (anonymous):

lol i know :P. i know them all

OpenStudy (anonymous):

my prof . dont alow anything except calculator :S

OpenStudy (anonymous):

lol funny. we weren't allowed a calculator on our exam for this class

OpenStudy (anonymous):

bigget "calculation" is factorials lol

OpenStudy (anonymous):

L(cos8t-sin8t )

OpenStudy (anonymous):

for laplace u dont need calculator we had it bec we had numerical math also and some other stuff xD jeff that is also table transform

OpenStudy (anonymous):

try and use this table and let us know if you still need help: http://mathworld.wolfram.com/LaplaceTransform.html

OpenStudy (anonymous):

lol we had an entire class for just this :P

OpenStudy (anonymous):

0.0 weird xD

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait 4 me.. i have 10 more qstion to solve.after i did it.i will up n share it.if wrong please correct 4 me

OpenStudy (anonymous):

if like this how to solve it (cos8t-sin8t)

OpenStudy (anonymous):

\[\frac{ s }{ s^2 + 8^2 } - \frac{ 8 }{ s^2 + 8^2 }\]

OpenStudy (anonymous):

must use 2 formul?

OpenStudy (anonymous):

yeap one for sin(a) and other for cos(at)

OpenStudy (anonymous):

sin(at)

OpenStudy (anonymous):

after tht

OpenStudy (anonymous):

after get the answer euler gve is it any working process to show?

OpenStudy (anonymous):

that is the final answer unless you want to replace 8^2 with 64 u can do it but u cant do anything more

OpenStudy (anonymous):

just like that?

OpenStudy (anonymous):

he just used the formulas from the laplace table for cos(at) and sin(at) ..

OpenStudy (anonymous):

(t4 e 3t)

OpenStudy (anonymous):

i am sorry buddy i totaly have to go , maybe euler can help you with the rest ^^ , good luck ;D

OpenStudy (anonymous):

ok tq

OpenStudy (anonymous):

let's denote the L[f(t)] as F(s) when you have e^(at)*f(t) the L[e^(at)*f(t)] = F(s-a) so you do the laplapce xform of f(t) then plug (s-a) for every s in F(s) \[L[t^4] = \frac{ 24 }{ s^5 }\] \[L[e^{3t}t^4] = \frac{ 24 }{ (s-3)^5 }\]